OpenStudy (anonymous):

Can someone walk me threw this problem? (Intro to Technical Mathematics) I have to do to the unit analysis to make the the necessary conservation. Change 88 feet per second to miles per hour.

OpenStudy (noelgreco):

\[\frac{ 88 feet }{1 \sec } \times \frac{ ?\sec}{1 hour }\times \frac{1 mile}{ ? feet }\] fill in the question marks and the units divide out to give you mph

OpenStudy (anonymous):

But how did you know to put it in that form?

OpenStudy (anonymous):

I get the first one. But after that I am lost...

OpenStudy (anonymous):

Still there?

OpenStudy (noelgreco):

Yes. This is called the factor-label method. 1. Write down the first quantity, in this case 88fps 2. Now choose either the time or the distance to work on first. I chose time. I put the unit (sec) that I want to change in the numerator of a fraction and put the time I want (hours) in the denominator. The reason I do this is because I want the sec to divide out (one in the numerator, and one in the denominator.) You're used to doing this with numbers, but here we're dividing out labels. 3. Put 3600 in front of the sec. Now your denominator is 1 hour, the numerator is also one hour, expressed in different units. When the numerator and the denominator of a fraction are equal, the value of the fraction is 1. When you multiply be one, you don't change the value of the original quantity, just the label (second to hours) 4. Do the same thing for all other units needing change.

OpenStudy (noelgreco):

Are you still trying? This is very useful to know once you learn it.